# Slender Trees and the Approximation Property

### v1

@@ Abstract.
We obtain a relatively simple criterion for when a forcing has the ${<}\,\delta$-approximation property, generalizing a result of Unger. Afterwards we apply this criterion to construct variants of Mitchell Forcing in order to answer questions posed by Mohammadpour.
@@ 1. Introduction
In his PhD thesis, Christoph Weiss introduced a two-cardinal generalisation of a $\kappa$-tree called a $(\kappa,\lambda)$-list along with two “thinness“ properties, namely thin and slender. The corresponding cardinal properties, stating that every “not too wide“ list has an ineffable branch, behave very differently in practice. The ineffable thin property, denoted $\operatorname{ITP}$, is consistent with e.g. the approachability property and forced by many orders which were originally conceived with the tree property in mind. On the other hand, the ineffable slender property, denoted $\operatorname{ISP}$, forces that $\operatorname{AP}$ fails and needs different preservation theorems compared to the tree property. The proof structures for consistency proofs involving $\operatorname{ISP}$ and $\operatorname{ITP}$ are the same: Given a list in the extension $V[G]$, lift some ground-model embedding $V\to M$ to an embedding $V[G]\to M[H]$ inside $V[H]$, thus obtaining an ineffable branch in $V[H]$ and “pull it back“ into $V[G]$ using the thinness assumption and properties of the extension $V[H]$ over $V[G]$. For thin lists, this is done using two branch lemmas stating that sufficiently closed forcings and forcings whose squares have small antichains do not add cofinal branches to thin lists. For $\delta$-slender lists however, we need that the pair $(V[G],V[H])$ has the so-called ${<}\,\delta$-approximation property. In practice, this property is much harder to establish and so many things concerning $\operatorname{ISP}$ are not known. In this paper, we obtain two important tools to prove consistency results involving $\operatorname{ISP}$: We obtain a general criterion for when a forcing order preserves the statement $\operatorname{ISP}$ and, more importantly, prove the following statement concerning the approximation property, generalizing a result of Unger from [13]:
@@
Let (×$(\mathbb{P}\times\mathbb{Q},R)$ be an iteration-like partial order and $\delta$ a cardinal. Assume that the base ordering $(\mathbb{P},b(R))$ is square-$\delta$-cc. and the term ordering (×$(\mathbb{P}\times\mathbb{Q},t(R))$ is strongly ${<}\,\delta$-distributive. Then (×$(\mathbb{P}\times\mathbb{Q},R)$ has the ${<}\,\delta$-approximation property.
Using this Theorem, we construct a number of variants of Mitchell Forcing to answer questions which were raised by Mohammadpour in [11]:
@@
    $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ holds and is indestructible under ${<}\,\kappa$-directed closed forcing.
@@ 2. Preliminaries
Our notation is fairly standard. We follow the convention that filters are upward closed, so that $q\leq p$ means that $q$ forces at least as much as $p$. $\mathbb{P}\upharpoonright p$ is defined as $\{q\in\mathbb{P}\;|\;q\leq p\}$. $V[\mathbb{P}]$ denotes an arbitrary extension by $\mathbb{P}$, i.e. “$V[\mathbb{P}]$ has property $P$“ means that for every $\mathbb{P}$-generic $G$, $V[G]$ has property $P$. Lastly, for any forcing order $\mathbb{P}$, we let $\Gamma_{\mathbb{P}}$ (omitting $\mathbb{P}$ when it is clear which order we are talking about) be the canonical $\mathbb{P}$-name such that $\Gamma_{\mathbb{P}}^{G}=G$ whenever $G$ is a $\mathbb{P}$-generic filter.
@@
Let $V\subseteq W$ be models of set theory with the same ordinals and $\kappa$ a regular cardinal in $V$. $(V,W)$ has the ${<}\,\kappa$-covering property if for any set $x\in W$, $x\subseteq V$ of size ${<}\,\kappa$ there is $y\in V$ of size ${<}\,\kappa$ such that $x\subseteq y$. A forcing order $\mathbb{P}$ has the $\kappa$-covering property if $1_{\mathbb{P}}$ forces that $(V,V[\mathbb{P}])$ has it.
If $(V,W)$ has the ${<}\,\kappa$-covering property, where $\kappa$ is a regular cardinal, $\kappa$ remains a regular cardinal in $W$: Otherwise there exists a cofinal subset of $\kappa$ in $W$ of size ${<}\,\kappa$. This set is covered by an element of $V$ of size ${<}\,\kappa$, contradicting the regularity of $\kappa$. If $\mathbb{P}$ is $\kappa$-cc. and $\dot{f}$ any $\mathbb{P}$-name for a function from an ordinal into $V$, then there are fewer than $\kappa$ possibilities for any value $\dot{f}(\check{\alpha})$, so we obtain:
@@
If $\mathbb{P}$ has the $\kappa$-cc., $\mathbb{P}$ has the $\kappa$-covering property.
@@
For completeness, we state the Product Lemma, which states that forcing with a product of orders can be viewed as successive forcing.
Lemma 2.6 (Product Lemma).
Let $\mathbb{P}$ and $\mathbb{Q}$ be notions of forcing. For $G\subseteq\mathbb{P}$, $H\subseteq\mathbb{Q}$, the following are equivalent:
  1. (1)
    $G\times H$ is $\mathbb{P}\times\mathbb{Q}$-generic over $V$.
  2. (2)
    $G$ is $\mathbb{P}$-generic over $V$ and $H$ is $\mathbb{Q}$-generic over $V[G]$.
  3. (3)
    $H$ is $\mathbb{Q}$-generic over $V$ and $G$ is $\mathbb{P}$-generic over $V[H]$.
The approximation property was introduced by Hamkins to show certain impossibility results concerning the indestructibility of large cardinal notions. However, it is also highly related to slender lists.
@@
Let $V\subseteq W$ be models of set theory and $\delta$ a cardinal in $V$. $(V,W)$ has the ${<}\,\delta$-approximation property if whenever $x\in W$ is such that $x\cap z\in V$ for every $z\in[x]^{<\delta}\cap V$, $x\in V$.
In his paper introducing the approximation property, Hamkins also showed that iterations $*\mathbb{P}*\mathbb{Q}$ where $\mathbb{P}$ is of size $<\delta$ and $\mathbb{Q}$ is ${<}\,\delta$-strategically closed have the ${<}\,\delta$-approximation property. Over the years, this criterion has been improved step by step (and criteria of a very different flavor such as strong properness have been introduced). In this paper, we will obtain a further weakening of the necessary assumptions.
Clubs and Stationarity in $[\lambda]^{<\kappa}$
Definition 2.8.
Let $\kappa\leq\lambda$ be cardinals of uncountable cofinality. $[\lambda]^{<\kappa}$ is defined as the set of all $x\subseteq\lambda$ with size ${<}\,\kappa$.
  1. (1)
    $C\subseteq[\lambda]^{<\kappa}$ is club in $[\lambda]^{<\kappa}$ if it is unbounded and closed under ascending unions.
  2. (2)
    $S\subseteq[\lambda]^{<\kappa}$ is stationary in $[\lambda]^{<\kappa}$ if it intersects every club.
The space $[\lambda]^{<\kappa}$ behaves much like $\kappa$: Namely, the intersection of fewer than $\operatorname{cf}(\kappa)$ club sets is club and the club filter is even closed under diagonal intersections, i.e. if $C_{\alpha}$ is club for every $\alpha<\lambda$, so is
$\Delta_{\alpha<\lambda}C_{\alpha}:=\{x\in[\lambda]^{<\kappa}\;|\;x\in\bigcap_{\alpha\in x}C_{\alpha}\}$
There is also a connection between clubs in $[\lambda]^{<\kappa}$ and functions $F\colon[\lambda]^{<\omega}\to[\lambda]^{<\kappa}$. If $F\colon[\lambda]^{<\omega}\to[\lambda]^{<\kappa}$, the set
$C_{F}:=\{x\in[\lambda]^{<\kappa}\;|\;\forall y\in[x]^{<\omega}\;F(y)\subseteq x\}$
is club in $[\lambda]^{<\kappa}$. It turns out that these sets generate the club filter (for a proof, see [10]):
Lemma 2.9.
Let $C$ be club in $[\lambda]^{<\kappa}$. Then there is $F\colon[\lambda]^{<\omega}\to[\lambda]^{<\kappa}$ such that $C_{F}\subseteq C$.
As a corollary, we obtain that clubs in $[\lambda]^{<\kappa}$ project to sets containing clubs in $[X]^{<\kappa}$ whenever $X\subseteq\lambda$:
Corollary 2.10.
Let $C\subseteq[\lambda]^{<\kappa}$ be club and $X\subseteq\lambda$ of size $\geq\kappa$. Then
$C\upharpoonright X:=\{N\cap X\;|\;N\in C\}$
contains a club in $[X]^{<\kappa}$.
Lists
@@ A $(\kappa,\lambda)$-list $f$ is thin if for every $x\in[\lambda]^{<\kappa}$,
    |$|\{f(z)\cap x\;|\;x\subseteq z\in[\lambda]^{<\kappa}\}|\;<\kappa$
@@
    is stationary in $[\lambda]^{<\kappa}$.
@@ 3. Strong Distributivity and Orders on Products
We expand the framework for working with arbitrary orders on products of sets by a new property and prove a result on when a forcing has the ${<}\,\delta$-approximation property. We first review definitions and results from [4].
@@
Let $\mathbb{P}$ be a poset and $\delta$ a cardinal. $\mathbb{P}$ is strongly ${<}\,\delta$-distributive if for every $p\in\mathbb{P}$ and all sequences $(D_{\alpha})_{\alpha<\delta}$ of open dense subsets of $\mathbb{P}\upharpoonright p$ there is a descending sequence $(p_{\alpha})_{\alpha<\delta}$ such that for every $\alpha<\delta$, $p_{\alpha}\in D_{\alpha}$.
@@
Let $\mathbb{P}$ be a forcing order, $\delta$ an ordinal. The completeness game $G(\mathbb{P},\delta)$ on $\mathbb{P}$ with length $\delta$ has players COM (complete) and INC (incomplete) playing elements of $\mathbb{P}$ with COM playing at even ordinals (and limits) and INC playing at odd ordinals. COM starts by playing $1_{\mathbb{P}}$, afterwards $p_{\alpha}$ has to be a lower bound of $(p_{\beta})_{\beta<\alpha}$. INC wins if COM is unable to play at some point $<\delta$. Otherwise, COM wins.
A forcing order $\mathbb{P}$ is ${<}\,\kappa$-distributive if and only if for all $\alpha<\kappa$, INC does not have a winning strategy in $G(\mathbb{P},\alpha)$. The proof adapts to show the following:
@@
$\mathbb{P}$ is strongly ${<}\,\kappa$-distributive if and only if INC does not have a winning strategy in $G(\mathbb{P},\kappa)$.
@@
A poset $\mathbb{P}$ is $\kappa$-strategically closed if COM has a winning strategy in $G(\mathbb{P},\kappa)$. It is ${<}\,\kappa$-strategically closed if for all $\alpha<\kappa$, COM has a winning strategy in $G(\mathbb{P},\alpha)$.
It is clear that if $\mathbb{P}$ is $\kappa$-strategically closed it is strongly ${<}\,\kappa$-distributive. However, there is no provable connection between ${<}\,\kappa$-strategic closure and strong ${<}\,\kappa$-distributivity.
@@
As variants of Mitchell Forcing are neither exactly products nor iterations, it is helpful to think of them as abstract orders on products of sets with certain regularity properties. This extends material from [4].
@@
Let $\mathbb{P}$ and $\mathbb{Q}$ be sets and $R$ a partial order on $\mathbb{P}\times\mathbb{Q}$. We will only consider orderings where for all $p,p^{\prime}$
$\exists q_{0},q_{1}((p,q_{0})R(p^{\prime},q_{1}))\longrightarrow\forall q((p,q)R(p^{\prime},q))$
If we want to reference this property, we will say that $\mathbb{P}\times\mathbb{Q}$ is based. We define the following partial orders:
@@
    (×$(\mathbb{P}\times\mathbb{Q},R)$ has the projection property if whenever $(p^{\prime},q^{\prime})R(p,q)$, there is $q^{\prime\prime}$ such that $(p,q^{\prime\prime})R(p,q)$ and $(p^{\prime},q^{\prime\prime})R(p^{\prime},q^{\prime})R(p^{\prime},q^{\prime\prime})$.
@@
    (×$(\mathbb{P}\times\mathbb{Q},R)$ has the refinement property if $p^{\prime}(b(R))p$ implies that $s(R,p^{\prime})$ refines $s(R,p)$, i.e. whenever $(p,q^{\prime})R(p,q)$ and $p^{\prime}(b(R))p$, also $(p^{\prime},q^{\prime})R(p^{\prime},q)$.
@@
    (×$(\mathbb{P}\times\mathbb{Q},R)$ has the mixing property if whenever $(p,q_{0}),(p,q_{1})R(p,q)$, there are $p_{0},p_{1}\in\mathbb{P}$ and $q^{\prime}$ with $(p,q^{\prime})\leq(p,q)$ such that $(p_{i},q^{\prime})\leq(p,q_{i})$.
We say that (×$(\mathbb{P}\times\mathbb{Q},R)$ is iteration-like if (×$(\mathbb{P}\times\mathbb{Q},R)$ is based and has the projection property, the refinement property and the mixing property.
The projection and refinement property hold in almost all cases, and always for iterations and products. They are necessary for most of the relevant techniques. The mixing property roughly states that we can mix elements of $\mathbb{Q}$ modulo $\mathbb{P}$ and holds e.g. in iterations $\mathbb{P}*\dot{\mathbb{Q}}$ if $\mathbb{P}$ is atomless. We also note that the term ordering is the disjoint union of the section orderings, so if the term ordering is ${<}\,\delta$-closed (strategically closed, strongly distributive etc.)for some $\delta$, so are all the section orderings (and vice versa).
@@
If (×$(\mathbb{P}\times\mathbb{Q},R)$ is based and has the projection and refinement property, the identity is a projection from $(\mathbb{P},b(R))\times(\mathbb{Q},s(R,1_{\mathbb{P}}))$ onto (×$(\mathbb{P}\times\mathbb{Q},R)$.
If (×$(\mathbb{P}\times\mathbb{Q},R)$ is an ordering on a product, forcing with $\mathbb{P}\times\mathbb{Q}$ can also be regarded as first forcing with $\mathbb{P}$ and then with a special ordering on $\mathbb{Q}$. However, this will not be of use to us in this paper. In [13], Unger showed that an iteration $*\mathbb{P}*\mathbb{Q}$, where $\mathbb{P}$ is square-$\delta$-cc. and $\mathbb{Q}$ is forced to be ${<}\,\delta$-closed, has the ${<}\,\delta$-approximation property. We will improve this result in two ways: Firstly, we work with an arbitrary order on a product, allowing us to apply the result more easily to variants of Mitchell Forcing and secondly, we only need a “distributivity“ assumption as opposed to one about closure. This is important, as, unlike ${<}\,\delta$-closure, strong ${<}\,\delta$-distributivity is preserved by forcing with $\delta$-cc. forcings.
@@
Let (×$(\mathbb{P}\times\mathbb{Q},R)$ be an iteration-like partial order and $\delta$ a cardinal. Assume $(\mathbb{P},b(R))^{2}$ is $\delta$-cc. and (×$(\mathbb{P}\times\mathbb{Q},t(R))$ is strongly $<\delta$-distributive. Then (×$(\mathbb{P}\times\mathbb{Q},R)$ has the $<\delta$-approximation property.
@@
Let (×$(\mathbb{P}\times\mathbb{Q},R)$ be an iteration-like partial order. If $(p,q)$ forces $\dot{x}\in V$ but not $\dot{x}=\check{y}$ for any $y\in V$, there are $q^{\prime\prime}\in\mathbb{Q}$, $p_{0},p_{1}b(R)p$ and $y_{0}\neq y_{1}$ such that $(p,q^{\prime\prime})R(p,q)$ and for $i\in 2$, $(p_{i},q^{\prime\prime})\Vdash\dot{x}=\check{y}_{i}$.

### v2

@@ Abstract.
We obtain a relatively simple criterion for when a forcing has the $\delta-$approximation property, generalizing a result of Unger. Afterwards we apply this criterion to construct variants of Mitchell Forcing in order to answer questions posed by Mohammadpour.
@@ 1. Introduction
In his PhD thesis, Christoph Weiss introduced a two-cardinal generalisation of a $\kappa$-tree called a $(\kappa,\lambda)$-list along with two “thinness“ properties, namely thin and slender. The corresponding cardinal properties, stating that every “not too wide“ list has an ineffable branch, behave very differently in practice. The ineffable tree property, denoted $\operatorname{ITP}$, is consistent with e.g. the approachability property and forced by many orders which were originally conceived with the tree property in mind. On the other hand, the ineffable slender tree property, denoted $\operatorname{ISP}$, forces that $\operatorname{AP}$ fails and needs different preservation theorems compared to the tree property.
The proof structures for consistency proofs involving $\operatorname{ISP}$ and $\operatorname{ITP}$ are the same: Given a list in the extension $V[G]$, lift some ground-model embedding $V\to M$ to an embedding $V[G]\to M[H]$ inside $V[H]$, thus obtaining an ineffable branch in $V[H]$ and “pull it back“ into $V[G]$ using the thinness assumption and properties of the extension $V[H]$ over $V[G]$. For thin lists, this is done using two branch lemmas stating that sufficiently closed forcings and forcings whose squares have small antichains do not add cofinal branches to thin lists. For $\delta$-slender lists however, we need that the pair $(V[G],V[H])$ has the so-called $\delta-$approximation property. In practice, this property is much harder to establish and so many results which have been shown for $\operatorname{TP}$ and $\operatorname{ITP}$ have not yet been established for $\operatorname{ISP}$. E.g. it was shown by Cummings and Foreman in [3] that the tree property can consistently hold at every $\aleph_{n+2}$, which was improved by Fontanella in [4] to show that $\operatorname{ITP}$ can consistently hold at every $\aleph_{n+2}$. On the other hand the only comparable result for $\operatorname{ISP}$ was obtained by Mohammadpour and Velickovic in [18] who showed using side condition forcing that consistently $\operatorname{ISP}(\omega_{2})$ and $\operatorname{ISP}(\omega_{3})$ can hold simultaneously.
In this paper, we obtain two important tools to prove consistency results involving $\operatorname{ISP}$: We obtain a general criterion for when a forcing order preserves the statement $\operatorname{ISP}$ and, more importantly, prove the following statement concerning the approximation property, generalizing a result of Unger from [20]:
@@
Let $(\mathbb{P}\times\mathbb{Q},R)$ be an iteration-like partial order and $\delta$ a cardinal. Assume that the base ordering $(\mathbb{P},b(R))$ is square-$\delta$-cc. and the term ordering $(\mathbb{P}\times\mathbb{Q},t(R))$ is strongly $\delta-$distributive. Then $(\mathbb{P}\times\mathbb{Q},R)$ has the $\delta-$approximation property.
We then demonstrate the robustness of the framework by constructing several new variants of Mitchell forcing in order to answer questions posed by Mohammadpour in a recent survey (see [17]). Recall that a cardinal $\kappa$ is $\lambda$-ineffable if $\kappa$ is inaccessible and $\operatorname{ISP}(\kappa,\lambda)$ holds.
@@
    $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ holds and is indestructible under $\kappa-$directed closed forcing.
@@
Acknowledgements.
This paper forms a part of the author’s PhD thesis. The author would like to thank his advisor, Heike Mildenberger, for many fruitful discussions surrounding this material. He would also like to thank Maxwell Levine for alerting him to the open questions guiding this paper and discussing them with him. Lastly, he would like the thank the referee for their thorough report and many helpful comments.
@@ 2. Preliminaries
Our notation is fairly standard. We follow the convention that filters are upward closed, so that $q\leq p$ means that $q$ forces at least as much as $p$. $\mathbb{P}\upharpoonright p$ is defined as $\{q\in\mathbb{P}\;|\;q\leq p\}$. $V[\mathbb{P}]$ denotes an arbitrary extension by $\mathbb{P}$, i.e. $V[\mathbb{P}]$ has property PP“ means that for every $\mathbb{P}$-generic $G$, $V[G]$ has property $P$. Lastly, for any forcing order $\mathbb{P}$, we let $\Gamma_{\mathbb{P}}$ (omitting $\mathbb{P}$ when it is clear which order we are talking about) be the canonical $\mathbb{P}$-name such that $\Gamma_{\mathbb{P}}^{G}=G$ whenever $G$ is a $\mathbb{P}$-generic filter.
@@
Let $V\subseteq W$ be models of set theory with the same ordinals and $\kappa$ a regular cardinal in $V$. $(V,W)$ has the $\kappa-$covering property if for any set $x\in W$, $x\subseteq V$ of size $\kappa$ in $W$ there is $y\in V$ of size $\kappa$ in $V$ such that $x\subseteq y$. A forcing order $\mathbb{P}$ has the $\kappa$-covering property if $1_{\mathbb{P}}$ forces that $(V,V[\mathbb{P}])$ has it.
If $(V,W)$ has the $\kappa-$covering property, where $\kappa$ is a regular cardinal, $\kappa$ remains a regular cardinal in $W$: Otherwise there exists a cofinal subset of $\kappa$ in $W$ of size $\kappa$. This set is covered by an element of $V$ of size $\kappa$, contradicting the regularity of $\kappa$. If $\mathbb{P}$ is $\kappa$-cc. and $\dot{f}$ any $\mathbb{P}$-name for a function from an ordinal into $V$, then there are fewer than $\kappa$ possibilities for any value $\dot{f}(\check{\alpha})$, so we obtain:
@@
If $\mathbb{P}$ has the $\kappa$-$c$., $\mathbb{P}$ has the $\kappa$-covering property.
@@
The approximation property was used implicitly by Mitchell in his proof that the tree property can consistently hold at successor cardinals. Later it was defined explicitly by Hamkins in [5]:
@@
Let $V\subseteq W$ be models of set theory and $\delta$ a cardinal in $V$. $(V,W)$ has the $\delta-$approximation property if $x\in V$ holds whenever $x\in W$ is such that $x\cap z\in V$ for every $z\in([x]^{<\delta})^{V}$. A poset $\mathbb{P}$ has the $\delta-$approximation property if $(V,V[G])$ has it for every $\mathbb{P}$-generic filter $G$.
In his paper introducing the approximation property, Hamkins also showed that iterations $\mathbb{P}*\mathbb{Q}$ where $\mathbb{P}$ is of size $\delta$ and $\mathbb{Q}$ is $\delta-$strategically closed have the $\delta-$approximation property. Over the years, this criterion has been improved step by step (and criteria of a very different flavor such as strong properness have been introduced). In this paper, we will obtain a further weakening of the necessary assumptions.
@@ A $(\kappa,\lambda)$-list $f$ is thin if for every $x\in[\lambda]^{<\kappa}$,
    $|\{f(z)\cap x\;|\;x\subseteq z\in[\lambda]^{<\kappa}\}|\;<\kappa$
@@
    is stationary in $\lambda]^{<\kappa}$.
@@ 3. Strong Distributivity and Orders on Products
We expand the framework for working with arbitrary orders on products of sets by a new property and prove a result on when a forcing has the $\delta-$approximation property.
We first review definitions and results from [7].
@@
Let $\mathbb{P}$ be a poset and $\delta$ a cardinal. $\mathbb{P}$ is strongly $\delta-$distributive if for every $p\in\mathbb{P}$ and all sequences $(D_{\alpha})_{\alpha<\delta}$ of open dense subsets of $\mathbb{P}\upharpoonright p$ there is a descending sequence $(p_{\alpha})_{\alpha<\delta}$ such that for every $\alpha<\delta$, $p_{\alpha}\in D_{\alpha}$.
@@
Let $\mathbb{P}$ be a forcing order, $\delta$ an ordinal. The completeness game $G(\mathbb{P},\delta)$ on $\mathbb{P}$ with length $\delta$ has players COM (complete) and INC (incomplete) playing elements of $\mathbb{P}$ with COM playing at even ordinals (and limits) and INC playing at odd ordinals. COM starts by playing $1_{\mathbb{P}}$, afterwards $p_{\alpha}$ has to be a lower bound of $(p_{\beta})_{\beta<\alpha}$. INC wins if COM is unable to play at some point $\delta$. Otherwise, COM wins.
A forcing order $\mathbb{P}$ is $\kappa-$distributive if and only if for all $\alpha<\kappa$, INC does not have a winning strategy in $G(\mathbb{P},\alpha)$. The proof adapts to show the following:
@@
$\mathbb{P}$ is strongly $\kappa-$distributive if and only if INC does not have a winning strategy in $G(\mathbb{P},\kappa)$.
@@
A poset $\mathbb{P}$ is $\kappa$-strategically closed if COM has a winning strategy in $G(\mathbb{P},\kappa)$. It is $\kappa-$strategically closed if for all $\alpha<\kappa$, COM has a winning strategy in $G(\mathbb{P},\alpha)$.
It is clear that if $\mathbb{P}$ is $\kappa$-strategically closed it is strongly $\kappa-$distributive. However, there is no provable connection between $\kappa-$strategic closure and strong $\kappa-$distributivity (see Example 3.11 and Example 3.12 in [7]).
@@
As variants of Mitchell Forcing are neither exactly products nor iterations, it is helpful to think of them as abstract orders on products of sets with certain regularity properties. We note that we will not assume that our partial orders are antisymmetric, i.e. for a partial order $\mathbb{P}$ there can be $p,p^{\prime}\in\mathbb{P}$ such that $p\leq p^{\prime}\leq p$ but $p\neq p^{\prime}$. This situation often occurs in iterated forcing, since, if $\tau$ and $\tau^{\prime}$ are different $\mathbb{P}$-names but $p\Vdash\tau=\tau^{\prime}$, we have $(p,\tau)\neq(p,\tau^{\prime})$ but $(p,\tau)\leq(p,\tau^{\prime})\leq(p,\tau)$.
@@
Let $\mathbb{P}$ and $\mathbb{Q}$ be nonempty sets and $R$ a partial order on $\mathbb{P}\times\mathbb{Q}$. We define the following related partial orders:
@@
  1. (i)
    $(\mathbb{P}\times\mathbb{Q},R)$ is based if for all $p,p^{\prime}$,
    $\exists q_{0},q_{1}((p,q_{0})R(p^{\prime},q_{1})\longrightarrow\forall q((p,q)R(p^{\prime},q)))$
  2. (ii)
    $(\mathbb{P}\times\mathbb{Q},R)$ has the projection property if whenever $(p^{\prime},q^{\prime})R(p,q)$, there is $q^{\prime\prime}\in\mathbb{Q}$ such that $(p,q^{\prime\prime})R(p,q)$ and $(p^{\prime},q^{\prime\prime})R(p^{\prime},q^{\prime})R(p^{\prime},q^{\prime\prime})$.
  3. (iii)
    $(\mathbb{P}\times\mathbb{Q},R)$ has the refinement property if $p^{\prime}(b(R))p$ implies that $s(R,p^{\prime})$ refines $s(R,p)$, i.e. whenever $(p,q^{\prime})R(p,q)$ and $p^{\prime}(b(R))p$, also $(p^{\prime},q^{\prime})R(p^{\prime},q)$.
  4. (iv)
    $(\mathbb{P}\times\mathbb{Q},R)$ has the mixing property if whenever $(p,q_{0}),(p,q_{1})R(p,q)$, there are $p_{0},p_{1}\in\mathbb{P}$ and $q^{\prime}\in\mathbb{Q}$ with $(p,q^{\prime})R(p,q)$ such that $(p_{i},q^{\prime})R(p,q_{i})$ for $i=0,1$.
We say that $(\mathbb{P}\times\mathbb{Q},R)$ is iteration-like if $(\mathbb{P}\times\mathbb{Q},R)$ is based and has the projection property, the refinement property and the mixing property.
The projection and refinement property hold in almost all cases, and always for iterations and products. They are necessary for most of the relevant techniques. The mixing property roughly states that we can mix elements of $\mathbb{Q}$ modulo $\mathbb{P}$ and holds e.g. in iterations $\mathbb{P}*\dot{\mathbb{Q}}$ if $\mathbb{P}$ is atomless:
Example 3.6.
[…]
@@
    The base ordering is simply equivalent to $(\mathbb{P},\leq_{\mathbb{P}})$ since $(p,\dot{q}_{0})R(p^{\prime},\dot{q}_{1})$ implies $p\leq_{\mathbb{P}}p^{\prime}$ and whenever $p\leq_{\mathbb{P}}p^{\prime}$, we can choose $\dot{q}_{0}=\dot{q}_{1}$.
@@
    The term ordering is equivalent to the usual notion of the term ordering on an iteration as defined by Krueger (see the discussion before [11, Proposition 2.1]).
@@
    For $p=1_{\mathbb{P}}$, the section ordering $s(R,p)$ is equal to the termspace forcing as defined by Laver (see e.g. [2, Chapter 22]).
Now let us verify that $(\mathbb{P}\times\tilde{\mathbb{Q}},R)$ has the defined properties:
  1. (i)
    $(\mathbb{P}\times\tilde{\mathbb{Q}},R)$ is based, since $(p,\dot{q}_{0})R(p^{\prime},\dot{q}_{1})$ implies $p\leq_{\mathbb{P}}p^{\prime}$ and thus that $(p,\dot{q})R(p^{\prime},\dot{q})$ for any $\dot{q}\in\tilde{\mathbb{Q}}$.
  2. (ii)
[…]
  3. (iii)
    $(\mathbb{P}\times\mathbb{Q},R)$ has the refinement property, since stronger conditions force more.
  4. (iv)
    $(\mathbb{P}\times\mathbb{Q},R)$ has the mixing property if $\mathbb{P}$ is atomless: Let $(p,\dot{q}_{0}),(p,\dot{q}_{1})R(p,\dot{q})$. Since $\mathbb{P}$ is atomless, we can find $p_{0},p_{1}\leq p$ which are incompatible. Using standard arguments on names, we can find $\dot{q}^{\prime}$ such that $p_{0}\Vdash\dot{q}^{\prime}=\dot{q}_{0}$ and $p_{1}\Vdash\dot{q}^{\prime}=\dot{q}_{1}$. Assume that conditions incompatible with both $p_{0}$ and $p_{1}$ force $\dot{q}^{\prime}=\dot{q}$. Then clearly $\dot{q}^{\prime}$ is as required.
Let us note that the term ordering is the disjoint union of the section orderings, so if the term ordering is $\delta-$closed (strategically closed, strongly distributive etc.) for some $\delta$, so are all the section orderings (and vice versa).
@@
If $(\mathbb{P}\times\mathbb{Q},R)$ is based and has the projection and refinement property, the identity is a projection from $(\mathbb{P},b(R))\times(\mathbb{Q},s(R,1_{\mathbb{P}}))$ onto $(\mathbb{P}\times\mathbb{Q},R)$.
Proof.
Denote by $R_{\pi}$ the order on $(\mathbb{P},b(R))\times(\mathbb{Q},s(R,1_{\mathbb{P}}))$. If $(p^{\prime},q^{\prime})R_{\pi}(p,q)$, $p^{\prime}(b(R))p$ and $(1,q^{\prime})R(1,q)$. By the refinement property, $(p,q^{\prime})R(p,q)$. By basedness $(p^{\prime},q^{\prime})R(p,q^{\prime})$. In summary,
$(p^{\prime},q^{\prime})R(p,q^{\prime})R(p,q)$
Assume $(p^{\prime},q^{\prime})R(p,q)$. This implies $p^{\prime}(b(R))p$. Furthermore, by basedness $(p,q)R(1_{\mathbb{P}},q)$. By the projection property there is $q^{\prime\prime}$ such that $(1_{\mathbb{P}},q^{\prime\prime})R(1_{\mathbb{P}},q)$ and $(p^{\prime},q^{\prime\prime})R(p^{\prime},q^{\prime})R(p^{\prime},q^{\prime\prime})$. So $(p^{\prime},q^{\prime\prime})R_{\pi}(p,q)$ and $(p^{\prime},q^{\prime\prime})R(p^{\prime},q^{\prime})$. ∎
If $(\mathbb{P}\times\mathbb{Q},R)$ is an ordering on a product, forcing with $\mathbb{P}\times\mathbb{Q}$ can also be regarded as first forcing with $\mathbb{P}$ and then with a special ordering on $\mathbb{Q}$. However, this will not be of use to us in this paper.
In [20], Unger showed that an iteration $\mathbb{P}*\mathbb{Q}$, where $\mathbb{P}$ is square-$\delta$-cc. and $\mathbb{Q}$ is forced to be $\delta-$closed, has the $\delta-$approximation property. We will improve this result in two ways: Firstly, we work with an arbitrary order on a product, allowing us to apply the result more easily to variants of Mitchell Forcing and secondly, we only need a “distributivity“ assumption as opposed to one about closure. This is important, as, unlike $\delta-$closure, strong $\delta-$distributivity is preserved by forcing with $\delta$-cc. forcings. So in particular, if we have a forcing $(\mathbb{P}\times\mathbb{Q},R)$ with its base ordering isomorphic to $\operatorname{Add}(\omega,X)$ for some set $X$ and a strongly $\omega_{1}-$distributive term ordering, it will satisfy the conclusion of Lemma 3.8 in any ccc. extension. This is a slight improvement of Lemma 5.4 in [22], where the square-ccc. is required instead.
@@
Let $(\mathbb{P}\times\mathbb{Q},R)$ be an iteration-like partial order and $\delta$ a cardinal. Assume $(\mathbb{P},b(R))^{2}$ is $\delta$-cc. and $(\mathbb{P}\times\mathbb{Q},t(R))$ is strongly $\delta-$distributive. Then $(\mathbb{P}\times\mathbb{Q},R)$ has the $\delta-$approximation property.
@@
Let $(\mathbb{P}\times\mathbb{Q},R)$ be an iteration-like partial order. If $(p,q)$ forces $\dot{x}\in V$ but for any $y\in V$, $(p,q)\not\Vdash\dot{x}=\check{y}$, there are $q^{\prime\prime}\in\mathbb{Q}$, $p_{0},p_{1}b(R)p$ and $y_{0}\neq y_{1}$ such that $(p,q^{\prime\prime})R(p,q)$ and for $i\in 2$, $(p_{i},q^{\prime\prime})\Vdash\dot{x}=\check{y}_{i}$.